OpenStudy (hi54321):

What is the slope-intercept form of the function that contains the points (6, 2) and (4, 8)? y = x + @kidrah69

OpenStudy (kidrah69):

well to find slope we need to use slope formula \[\frac{ y2-y1 }{ x2-x1 }\]

OpenStudy (kidrah69):

\[\frac{ 8-2 }{ 4-6 }\]

OpenStudy (kidrah69):

solve that

OpenStudy (irishboy123):

function that contains the points (6, 2) and (4, 8)? draw it always....

OpenStudy (hi54321):

first one to give me answer gets a medal

OpenStudy (kidrah69):

no

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ % (a,b) &({\color{red}{ 6}}\quad ,&{\color{blue}{ 2}})\quad % (c,d) &({\color{red}{ 4}}\quad ,&{\color{blue}{ 8}}) \end{array} \\\quad \\ % slope = m slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}} \\ \quad \\ % point-slope intercept y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values and solve for "y"}\\ \qquad \uparrow\\ \textit{point-slope form}\) get it, and you'd get a medal